OpenStudy (anonymous):

Solve 3the square root of x minus three) – 5 = 16 x = 5 x = 19 x = 52 x = 67

OpenStudy (cwrw238):

sqrt(x - 3) - 5 = 16 sqrt(x - 3) = 21 now square both sides

OpenStudy (anonymous):

how did you get rid of the cubed root? \[\sqrt[3]{x-3} - 5 = 16\]

OpenStudy (cwrw238):

oh - missed the 3 sorry

OpenStudy (cwrw238):

r u sure thats right not 3 times the square root of x - 3?

OpenStudy (anonymous):

it probably is three times the square root, im sorry

OpenStudy (anonymous):

so squared it would be 3(x-3) = 441?

OpenStudy (cwrw238):

lol ok so 3 sqrt(x - 3) = 21 sqt(x - 3) = 21/3 = 7 x- 3 = 49

OpenStudy (anonymous):

so x= 52?

OpenStudy (cwrw238):

yep

OpenStudy (anonymous):

cool, thank you!!

OpenStudy (cwrw238):

yw

OpenStudy (cwrw238):

the aim is to isolate the square root bit then square both sides